Use the square root property to solve the equation, show steps and explain
x^2=36
(x^2-36)=0 Factor out the left-side
(x-6) (x+6)=0 (? lost after this point)
(x-6)=0 or (x+6)=0
Thank you

Answer by checkley71(8403)

(X-6)(X+6)=0 THEN SETR EACH FACTOR TO 0 YOU GET X-6=0 OR ADDING 6 TO BOTH SIDES OF THE EQUATION YOU GET X-6+6=0+6 OR X=6 & X+6=0 OR SUBTRACTING 6 FROM BOTH SIDES YOU GET X+6-6=0-6 OR X=-6